the sum of the digits of a 2-digit number is 7. The tens digit is one less than 3 times the units digit. Find the number.
Answer by Mathtut(3670)

lets call our number ab....
:
a+b=7.........eq 1
a=3b-1.......eq 2
:
take a's value from eq 2 and plug it into eq 1
:
3b-1+b=7
:
4b=8
:
b=2
:
a=3(2)-1=5
:
so our number is 52